# Scaling Up On-Device LLMs via Active-Weight Swapping Between DRAM and Flash

### v1

@@ 7.4. Power and energy consumption
We also evaluate the power consumption and energy efficiency of ActiveFlow compared to the baseline on Device 1, as shown in Fig. 19. During the decoding process, ActiveFlow’s average power consumption is 1.86W, which is 27.34% lower than llama.cpp. This reduction is primarily due to computation wait time while fetching data from flash memory in our overlap pipeline. However, since ActiveFlow reduces the required weight to compute and load, the processing time per token is also reduced. As memory cost decreases, the energy consumption per token further declines. Specifically, at a 1.3GB memory cost, our system achieves a 53% energy reduction compared to llama.cpp.
@@ 8. Related Works
Sparsity in LLMs. Sparsity in LLMs have been the focus of many research efforts. Mirzedeh et al. (Mirzadeh et al., 2023) propose replacing the ReLU activation function in LLMs to reduce computation and weight transfer. HiRE (L et al., 2024) introduces high-recall approximate Top-K estimation for sparse LLMs. Prosparse (Song et al., 2024a) leverages the sparsity of the ReLU function and the gated branches within the feed-forward network (FFN) to predict model sparsity. Q-Sparse (Wang et al., 2024c) is an effective approach to training sparsely activated LLMs from scratch. TEAL (Liu et al., 2025a) is a training-free method that applies magnitude-based activation sparsity to LLM activations. Compared to these works, ActiveFlow further optimizes LLM weights through self-distillation, taking into account quantization and sparsity, and achieves SOTA scaling law on accuracy in both generation and downstream tasks.
Efficient LLM inference system. Based on the sparse activation characteristics of LLMs, previous work has designed systems for efficient model loading and inference. DeJavu (Liu et al., 2023) is a system that employs a low-cost algorithm to dynamically predict contextual sparsity. Alizadeh et al. (Alizadeh et al., 2024) propose a system that enables effective inference of LLMs on devices with limited memory. PowerInfer (Song et al., 2024b) is a CPU-GPU hybrid inference engine. ActiveFlow stands out from these works by (1) supporting any LLMs without the limitation of the ReLU activation function, and (2) providing elastic memory management for various devices with limited memory.
Static pruning techniques. Static pruning and quantization methods have been widely explored for compressing large language models. For instance, CFSP (Wang et al., 2025a) leverages structural redundancy for efficient pruning, while DB-LLM (Chen et al., 2024) and PB-LLM (Shang et al., 2023) propose dynamic block pruning strategies to reduce model parameters. In addition, RIA (Biswas et al., 2025) introduces regularization-based approaches to enhance quantization accuracy. These techniques enable substantial reductions in model size and computation without extensive retraining. However, they necessitate additional processing of model weights, making dynamic task adaptation and flexible loading challenging. ActiveFlow is nor only compatible with static weight optimize techniques but also supports dynamic processing.
Knowledge distillation for LLMs. Quantization-aware distillation has emerged as an effective paradigm to compress LLMs while mitigating accuracy loss. BitDistiller (Du et al., 2024) exemplifies this by integrating low-bit quantization with teacher-student distillation, achieving competitive performance under tight resource constraints. Other works (Zhao and Zhao, 2024) further refine the alignment between teacher and student models under quantization, balancing compression and task performance. Such approaches complement dynamic pruning methods and are especially attractive for deployment on devices with limited memory. ActiveFlow considers both quantization and sparsity factors during self-distillation, achieving SOTA in accuracy for both generation and downstream tasks.

### v2

@@ 7.4. Power and energy consumption
We evaluate power and energy efficiency of ActiveFlow on Device 1 (Fig. 20). ActiveFlow reduces average power consumption by 27.34% compared to llama.cpp due to reduced computation wait time in the overlap pipeline, and further lowers energy per token as memory cost decreases, achieving up to 53% reduction at 1.3GB memory usage.
@@ 8. Related Works
Sparsity in LLMs. Sparsity in LLMs has been the focus of many research efforts. Mirzadeh et al. (Mirzadeh et al., 2023) propose replacing the ReLU activation function in LLMs to reduce computation and weight transfer. HiRE (L et al., 2024) introduces high-recall approximate Top-K estimation. Prosparse (Song et al., 2024a) leverages the sparsity of ReLU and gated branches in FFNs to predict model sparsity. Q-Sparse (Wang et al., 2024c) trains sparse LLMs from scratch, while TEAL (Liu et al., 2025a) applies magnitude-based sparsity without retraining. However, these methods either depend on ReLU-based architectures or lack mechanisms to recover accuracy under high sparsity. InfiniGen(Lee et al., 2024b), FlexGen (Sheng et al., 2023) and related work primarily focus on KV cache optimization, KV cache dominants the LLM memory usage for long context scenarios (¿32K tokens). while we target weight memory optimization: usually determined by LLM’s weights.
Efficient LLM inference system. Several system-level efforts focus on exploiting sparsity for efficient inference. DejaVu (Liu et al., 2023) predicts contextual sparsity with lightweight algorithms, Alizadeh et al. (Alizadeh et al., 2024a) optimize inference on limited-memory devices, and PowerInfer (Song et al., 2024b) (and its extension PowerInfer-2) design CPU–GPU hybrid engines. These works primarily target ReLU-based models and FFN layers, often relying on heavy predictors (GB-level memory) to skip zero activations. Yet modern LLMs such as LLaMA and Mixtral adopt non-ReLU activations for accuracy (Touvron et al., 2023), limiting the applicability of these methods. LLM-in-Flash (Alizadeh et al., 2024b) streams weights from flash with fine-grained prefetching to reduce DRAM usage, but its efficiency is limited by flash bandwidth and latency, especially for compute-intensive layers.
Our distinction.ActiveFlow eliminates the ReLU dependency and predictor overhead by targeting all weights (Attention and FFN) in modern non-ReLU LLMs. It targets all weights (both Attention and FFN) and eliminates the need for predictors. It introduces (1) cross-layer active weight preloading, generalizing cross-layer similarity, and (2) sparsity-aware self-distillation to recover accuracy under high sparsity. Combined with an LFU-based cache driven by activation statistics, our approach consistently achieves higher hit rates (e.g., >>70% vs. $\sim$55%) and ensures strict memory budgets, enabling reliable edge deployment.
Static pruning techniques. Static pruning and quantization are established methods for compressing large language models (e.g., CFSP (Wang et al., 2025a), DB-LLM (Chen et al., 2024), and RIA (Biswas et al., 2025)). Although effective at reducing model size and computation, these static approaches require offline processing of model weights, which limits flexibility for dynamic tasks. Our system, ActiveFlow, is not only compatible with these static techniques but also uniquely supports dynamic processing to address this challenge.